# A Statistical Evaluation of Indoor LoRaWAN Environment-Aware Propagation for 6G: MLR, ANOVA, and Residual Distribution Analysis

### v2

@@ Abstract
Modeling path loss in indoor LoRaWAN technology deployments is inherently challenging due to structural obstructions, occupant density and activities, and fluctuating environmental conditions. This study proposes a two-stage approach to capture and analyze these complexities using an extensive dataset of 1,328,334 field measurements collected over 6 months in a single-floor office at the University of Siegen’s Hölderlinstraße Campus, Germany. First, we implement a multiple linear regression (MLR) model that includes the traditional propagation metrics (distance, structural walls) and an extension with proposed environmental variables (relative humidity, temperature, carbon dioxide (CO2), particulate matter, and barometric pressure). Using analysis of variance (ANOVA), we demonstrate that adding these environmental factors can reduce unexplained variance by 42.32%. Secondly, we examine residual distributions by fitting five candidate probability distributions: Normal, Skew-Normal, Cauchy, Student’s $t$, and Gaussian Mixture Models (GMM) with 1 to 5 components. Our results show that a four-component GMM captures the residual heterogeneity of indoor signal propagation most accurately, significantly outperforming single-distribution approaches. Given 6G’s push for ultra‐reliable, context‐aware communications, our analysis shows that environment‐aware modeling can substantially improve LoRaWAN network design in dynamic indoor IoT deployments.
@@
6G, ANOVA, environmental factors, indoor LoRaWAN, IoT, Multiple Linear Regression (MLR), path loss modeling, residual distribution analysis
@@ I Introduction
We propose a robust statistical approach designed to overcome the limitations prevalent in current indoor LoRaWAN propagation studies. Our objective is to clearly isolate and quantify the combined effects of distance, structural barriers, and dynamic environmental variables, providing deeper insights into their contributions to indoor signal attenuation. Leveraging a six-month empirical campaign comprising $1,328,334$ observations collected in an operational office environment, our analysis ensures adherence to statistical validity criteria essential for accurate modeling and practical network optimization. Precisely, our analytical framework consists of the following key components:
@@
    A multiple linear regression (MLR) model that captures distance, wall composition, and ambient environmental metrics of relative humidity, temperature, carbon dioxide (CO2), particulate matter (PM2.5) and barometric pressure,
@@
    A detailed residual distribution analysis to check compliance with theoretical assumptions on error normality and homoscedasticity.
In real-world indoor deployments, incorporating environmental data is crucial for improving path loss predictions and enabling power optimization, accurate localization, and robust network design. Although continuously monitoring these parameters may increase power consumption and require adaptive calibration to cope with dynamic occupancy and Heating, Ventilation, and Air Conditioning (HVAC)-induced variations, these measurements provide actionable insights. By deploying low-cost sensors alongside LoRaWAN nodes, practitioners can enable real-time path loss adjustments and design fade margins based on multimodal shadowing, thereby ensuring energy-efficient and resilient operation. Therefore, our study offers a statistically robust and practically viable toolkit for next-generation indoor LoRaWAN path loss modeling, paving the way for more resilient and energy-efficient deployments in the context of 6G and IoT applications.
@@ III-A Measurement Experimental Setup
In this study, we deployed a sensor network of $6$ LoRaWAN end devices (ED1 to ED6) as shown in Figure 1 on the eighth floor of an academic building at the University of Siegen’s Hölderlinstraße Campus in Germany, to systematically measure indoor path loss and ambient environmental parameters. We placed the end devices at distances ranging from $8\,\mathrm{m}$ to $40\,\mathrm{m}$ from a centrally located Wirnet iFemtoCell indoor gateway, as shown in Figure 2.
@@ IV-A MLR Model Results
IV-B Statistical Significance of Environmental Factors
ANOVA tests whether each factor or factor interaction explains a significant portion of the residual variance [23]. Predictor significance in the MLR model was evaluated through individual predictor $t$-tests and a Type II ANOVA to evaluate each variable’s unique contribution and combined effects within the model (see Table II). All factors except temperature demonstrated highly significant effects ($p<0.001$), with $t$-values ranging from $-13.59$ (pressure) to $-748.46$ (SNR), and corresponding $F$-statistics between $184.6$ and $560,185$. Temperature ($t=-11.06$, $F=1.35$, $p=0.245$) had limited marginal contribution, likely due to collinearity with other environmental parameters. Although not strictly environmental, SNR exhibited the strongest statistical significance due to its direct relationship to received signal quality. These results’ strong statistical and physical coherence highlights the necessity of integrating environmentally aware variables into indoor propagation models, with their direct impacts discussed in Section  IV-A.
@@ IV-B Statistical Significance of Environmental Factors
V Residual Distribution Analysis
@@ VI Conclusion and Future Directions
This work introduced an integrated approach for indoor LoRaWAN path loss modeling that combines multiple linear regression (MLR), ANOVA-based significance testing, and multimodal residual analysis. Utilizing a dataset of $1,328,334$ field measurements, we demonstrated that incorporating environmental factors (relative humidity, temperature, barometric pressure, particulate matter, and carbon dioxide) reduced unexplained variance by $42.3\%$. A key finding is that a four-component Gaussian Mixture Model (GMM) accurately reflects heterogeneous error patterns, likely driven by fluctuating occupancy or microclimatic conditions. Therefore, the proposed environment-aware strategy holds promise for more resilient indoor LoRaWAN deployments, aligning with the needs of 6G and IoT systems that demand consistently robust network performance and energy efficiency.

### v3

@@ Abstract
Modeling path loss in indoor LoRaWAN technology deployments is inherently challenging due to structural obstructions, occupant density and activities, and fluctuating environmental conditions. This study proposes a two-stage approach to capture and analyze these complexities using an extensive dataset of 1,328,334 field measurements collected over 6 months in a single-floor office at the University of Siegen’s Hölderlinstraße Campus, Germany. First, we implement a multiple linear regression (MLR) model that includes the traditional propagation metrics (distance, structural walls) and an extension with proposed environmental variables (relative humidity, temperature, carbon dioxide (CO2), particulate matter, and barometric pressure). Using analysis of variance (ANOVA), we demonstrate that adding these environmental factors can reduce unexplained variance by 42.32%. Secondly, we examine residual (shadow fading) distributions by fitting five candidate probability distributions: Normal, Skew-Normal, Cauchy, Student’s $t$, and Gaussian Mixture Models (GMM) with 2 to 5 components. Our results show that a four-component GMM captures the residual heterogeneity of indoor signal propagation most accurately, significantly outperforming single-distribution approaches. Given 6G’s push for ultra‐reliable, context‐aware communications, our analysis shows that environment‐aware modeling can substantially improve LoRaWAN network design in dynamic indoor IoT deployments.
@@
6G, ANOVA, environmental factors, indoor LoRaWAN, IoT, Multiple Linear Regression (MLR), path loss modeling, residual distribution analysis, shadow fading
@@ I Introduction
We propose a robust statistical approach designed to overcome the limitations prevalent in current indoor LoRaWAN propagation studies. Our objective is to clearly isolate and quantify the combined effects of distance, structural barriers, and dynamic environmental variables, providing deeper insights into their contributions to indoor signal attenuation. Leveraging a six-month empirical campaign comprising $1,328,334$ observations collected in an operational office environment, our analysis ensures adherence to statistical validity criteria essential for accurate modeling and practical network optimization. A comprehensive description of this dataset can be found in our previous work [6]. In this work, our analytical framework consists of the following key components:
@@
    A multiple linear regression (MLR) model that captures distance, wall composition, and ambient environmental metrics of relative humidity, temperature, carbon dioxide (CO2), particulate matter (PM2.5), and barometric pressure,
@@
    A detailed shadow fading (residual distribution) analysis to check compliance with theoretical assumptions on error normality, autocorrelation, and homoscedasticity.
In real-world indoor deployments, incorporating environmental data is crucial for improving path loss predictions [7], which enables power optimization, accurate localization, and robust network design. Although continuously monitoring these parameters may increase power consumption and require adaptive calibration to cope with dynamic occupancy and Heating, Ventilation, and Air Conditioning (HVAC)-induced variations, these measurements provide actionable insights. By deploying low-cost sensors alongside LoRaWAN nodes, practitioners can enable real-time path loss adjustments and design fade margins based on multimodal shadowing, thereby ensuring energy-efficient and resilient operation. Therefore, our study offers a statistically robust and practically viable toolkit for next-generation indoor LoRaWAN path loss modeling, paving the way for more resilient and energy-efficient deployments in the context of 6G and IoT applications.
@@ III-A Measurement Experimental Setup
In this study, we set up a sensor network of $6$ LoRaWAN end devices (ED1 to ED6), as shown in Figure 1. This was deployed on the eighth floor (approximately $240\,\mathrm{m}^{2}$, $290\,\mathrm{m}$ above sea level) of an academic building at the University of Siegen in Germany. The objective was to systematically measure indoor path loss and ambient environmental parameters. We placed the end devices at distances ranging from $8\,\mathrm{m}$ to $40\,\mathrm{m}$ from a centrally located Wirnet iFemtoCell indoor gateway, as shown in Figure 2.
@@ IV-A MLR Model Results
IV-B Statistical Significance of Environmental Factors
ANOVA tests whether each factor or factor interaction explains a significant portion of the residual variance [24]. Predictor significance in the MLR model was evaluated through individual predictor $t$-tests and a Type II ANOVA to evaluate each variable’s unique contribution and combined effects within the model (see Table II). All factors except temperature demonstrated highly significant effects ($p<0.001$), with $t$-values ranging from $-13.59$ (pressure) to $-748.46$ (SNR), and corresponding $F$-statistics between $184.6$ and $560,185$. Temperature ($t=-11.06$, $F=1.35$, $p=0.245$) had a limited marginal contribution, likely due to collinearity with other environmental parameters. Although not strictly environmental, SNR exhibited the strongest statistical significance due to its direct relationship to received signal quality. These results’ strong statistical and physical coherence highlight the necessity of integrating environmentally aware variables into indoor propagation models, with their direct impacts discussed in Section  IV-A.
V Shadow Fading / Residual Distribution Analysis
@@ VI Conclusion and Future Directions
This work introduced an integrated approach for indoor LoRaWAN path loss modeling that combines multiple linear regression (MLR), ANOVA-based significance testing, and multimodal residual (shadow fading) analysis. Utilizing a dataset of $1,328,334$ field measurements, we demonstrated that incorporating environmental factors (relative humidity, temperature, barometric pressure, particulate matter, and carbon dioxide) reduced unexplained variance by $42.3\%$. A key finding is that a four-component Gaussian Mixture Model (GMM) accurately reflects heterogeneous error patterns, likely driven by fluctuating occupancy or microclimatic conditions. Therefore, the proposed environment-aware strategy holds promise for more resilient indoor LoRaWAN deployments, aligning with the needs of 6G and IoT systems that demand consistently robust network performance and energy efficiency.
@@
Data Availability
The dataset and analysis scripts used for this work are openly available at: https://github.com/nahshonmokua/LoRaWAN-IndoorPathLossModeling-ML